OpenStudy (anonymous):

Find three points that solves the equation 2x-3y=-6

OpenStudy (anonymous):

EX:(3,2)

OpenStudy (texaschic101):

pick any number for x and we can solve this.....pick a number

OpenStudy (anonymous):

3

OpenStudy (texaschic101):

okay...now sub 3 in for x and find y 2x - 3y = -6 2(3) - 3y = -6 6 - 3y = -6 -3y = -6 - 6 -3y = -12 y = 4 so one of your points is (3,4)

OpenStudy (anonymous):

Ok need 2 more

OpenStudy (ciarán95):

The best option here is to pick a value for x in the equation above, substitute it in and then solve the remaining variable y. In this way you will get pairs of values (x = something, y = something) which satisfy the equation.

OpenStudy (texaschic101):

keep doing this...sub any number in for x and find y and you will have your points

OpenStudy (anonymous):

2

OpenStudy (texaschic101):

so sub 2 in for x and solve for y. Try it...I will stay and check it for you

OpenStudy (anonymous):

(2,2) ???

OpenStudy (texaschic101):

2x - 3y = -6 2(2) - 3y = -6 4 - 3y = -6 -3y = -6 - 4 -3y = -10 y = 10/3 some numbers will leave you with a fraction...but thats ok. another set of points is : (2,10/3)

OpenStudy (anonymous):

4

OpenStudy (texaschic101):

2(4) - 3y = -6 8 - 3y = -6 -3y = -6 - 8 -3y = - 14 y = 14/3 your points are : (4,14/3)

OpenStudy (texaschic101):

so we have : (3,4) (2,10/3) (4,14/3)